solve x
logx 25 = -2
***the x is subscript of log

Found 2 solutions by venugopalramana, AnlytcPhil:
Answer by venugopalramana(3286)

solve x
logx 25 = -2
***the x is subscript of log
X^(-2)=25
1/(X^2)=25
1/X=5
X=1/5

Answer by AnlytcPhil(1810)

solve for x 
logx 25 = -2 

Use the rule that says: 

A logarithmic equation of the form

  logBA = C

is equivalent to and can be rewritten
as the exponential equation

     BC = A

So we rewrite

   logx25 = -2

as

    x-2 = 25

Then we write the x-2 as 1/x2

   1/x2 = 25

Multiply both sides by x2:

      1 = 25x2

Divide both sides by 25 

   1/25 = x2 

Take positive square roots of both sides
(logarithm bases must be positive and ¹ 1,
by definition)
    
    1/5 = x

Edwin